OpenStudy (tester97):

what are the solutions to the quadratic equation of 2x^2-16x+32=0? and how would i solve this?

OpenStudy (anonymous):

Step 1: Factor out a 2, the GCF. I will wait till you do that.

OpenStudy (anonymous):

2x2 – 16x + 32 = 0 2(x^2-8x + 16) = 0 (x^2-8x+16) = 0 (x^2-4x -4x + 16) = 0 (x(x-4)-4(x-4)) = 0 (x-4)(x-4) = 0 (x-4)^2 = 0 (x-4)^2 = 0 x - 4 = 0 x = 4 Sorry, you don't need to quadratic formula to solve this.

OpenStudy (anonymous):

I have proof, I checked it on Wolframalpha http://www.wolframalpha.com/input/?i=solve%282x%5E2+%E2%80%93+16x+%2B+32+%3D+0%2C+x%29

OpenStudy (tester97):

omg thx buddy that explanation helped me alot! u have just earned a fan ^_^

OpenStudy (anonymous):

Anytime :) Glad I could assist you @tester97